OpenStudy (anonymous):

How would you graph the solution set -2(x-1) is less than or equal to 8?

jimthompson5910 (jim_thompson5910):

First you need to solve for x

jimthompson5910 (jim_thompson5910):

What do you get when you do this

jimthompson5910 (jim_thompson5910):

close, more like -2(x-1) <= 8 x-1 >= 8/(-2) x-1 >= -4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok I get it now thanks so you just have to graph it like that

jimthompson5910 (jim_thompson5910):

completely solve for x, then you can graph after that

OpenStudy (anonymous):

K

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

x >= -3

jimthompson5910 (jim_thompson5910):

good, so what does the graph look like?

OpenStudy (anonymous):

It would be like a dot on -3 with an arrow going left

jimthompson5910 (jim_thompson5910):

the arrow will go right, but you're correct: a closed circle is plotted on -3